# Bipartite and tripartite entanglement in pure dephasing relativistic spin-boson model

### v1

@@ B.2 Two gapless qubits in Minkowski spacetime
Define the distance $L$ between the detectors as $L\coloneqq|{\bm{x}}_{\mathrm{A,0}}-{\bm{x}}_{\mathrm{B,0}}|$. From Eq. (45), the reduced density matrix $\rho_{\mathrm{AB}}(t)$ in the basis {$\{\ket{++},\ket{+-},\ket{-+},\ket{--}\}$ reads
@@
×$\quad\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$ (50c)
@@
×$\quad\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$ (50d)
@@
Here, ${}_{0}\mathcal{F}_{1}$ is the regularized generalized hypergeometric function and $\Gamma(z)$ is the Gamma function [95, 96]. We note that the phase term
@@
as $\mathrm{Im}(\eta_{\bm{k}}(t)\bm{s}\cdot\bm{F_{k}},\eta_{\bm{k}}(t)\bm{r}\cdot\bm{F_{k}})_{\mathcal{H}}=0$ for a real smearing function $F^{(j)}$. [^11] This is consistent with the fact that the causal propagator $E(f,g)$ is zero when the supports of $f$ and $g$ are causally disconnected, as $\mathrm{Im}(f_{\bm{k}},g_{\bm{k}})_{\mathcal{H}}$ is the “${\bm{k}}$-space version” of $E(f,g)$. See [29].
B.3 Three gapless qubits in the equilateral triangular configuration
We now consider three static gapless detectors A, B, and C, in $(n+1)$-dimensional Minkowski spacetime. In this tripartite case, we have the freedom to choose the spatial configuration of the detectors. The simplest one is the equilateral triangle configuration [42, 28, 43], namely, each detector is located at a corner of an equilateral triangle at each time $t$. Such a symmetrical configuration simplifies the calculation drastically, as the proper distance between any pair of the detectors is the same $L$.
@@ B.3 Three gapless qubits in the equilateral triangular configuration
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,,$
@@
×$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,,$

### v3

@@ A.1 Two gapless qubits in Minkowski spacetime
[…]
@@
$\quad\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$, (50c)
@@
$\quad\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$, (50d)
@@
Here, $_{0}\mathcal{F}_{1}$ is the regularized generalized hypergeometric function and $\Gamma(z)$ is the Gamma function [NIST:DLMF, abramowitz1965handbook]. We note that the phase term
@@
as $\mathrm{Im}(\eta_{\bm{k}}(t)\bm{s}\cdot\bm{F_{k}},\eta_{\bm{k}}(t)\bm{r}\cdot\bm{F_{k}})_{\mathcal{H}}=0$ for a real smearing function $F^{(j)}$. [^11] This is consistent with the fact that the causal propagator $E(f,g)$ is zero when the supports of $f$ and $g$ are causally disconnected, as $\mathrm{Im}(f_{\bm{k}},g_{\bm{k}})_{\mathcal{H}}$ is the $\bm{k}$}-space version” of $E(f,g)$. See [Tjoa2024interacting].
A.2 Three gapless qubits in the equilateral triangular configuration
We now consider three static gapless detectors A, B, and C, in $(n+1)$-dimensional Minkowski spacetime. In this tripartite case, we have the freedom to choose the spatial configuration of the detectors. The simplest one is the equilateral triangle configuration [mendez2022tripartite, mendez2023tripartite, membrere2023tripartite], namely, each detector is located at a corner of an equilateral triangle at each time $t$. Such a symmetrical configuration simplifies the calculation drastically, as the proper distance between any pair of the detectors is the same $L$.
@@ A.2 Three gapless qubits in the equilateral triangular configuration
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))$
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{A})}\eta_{\bm{k}}^{*}(t)))\,$,
@@
$\times\omega_{0}(W(F_{\bm{k}}^{(\text{B})}\eta_{\bm{k}}^{*}(t)))\omega_{0}(W(F_{\bm{k}}^{(\text{C})}\eta_{\bm{k}}^{*}(t)))\,$,